Two charges, q1 and, q2, are located at the origin and at (0.50 m, 0), respectively. Where on the x-axis must a third charge, q3, of arbitrary sign be placed to be in electrostatic equilibrium if

(a) q 1 and q2 are like charges of equal magnitude.

(b) q1 and q2 are unlike charges of equal magnitude.

(c) q1=+3.0 μC and q2=−7.0 μC?


1
Expert's answer
2021-07-01T04:57:00-0400

a)


"x=0.25\ m"


b) Nowhere. It is not possible because one charge will always attract and the other repel q3 unequally.

c)


"\frac{3}{x^2}=\frac{7}{(x+0.5)^2}\\x=-0.95\ m\\y=0\ m"